# Knowledge Transfer from Simple to Complex: A Safe and Efficient Reinforcement Learning Framework for Autonomous Driving Decision-Making

### v3

@@ 6 Conclusion
This paper introduces a novel framework, S2CD, based on knowledge transfer techniques. The S2CD framework first trains a teacher model in a simplified simulation environment and then utilizes this model to guide the student agent in a more complex environment, enhancing the safety and efficiency of the training process. The paper also presents an innovative RL algorithm, ACPPO, which leverages samples generated by both teacher and student policies while dynamically adjusting the clipping factor based on sample importance, thereby improving learning efficiency. Moreover, the KL divergence between the teacher’s and student’s policies is incorporated as a constraint in model updates, solved using the Lagrangian method to enable the student agent to quickly adopt the teacher’s policy. To further enhance the student agent’s performance, a gradual weaning strategy reduces teacher intervention, ensuring that in the later stages of training, the student becomes more autonomous and explores optimal policies independently. Experimental results in highway lane-change scenarios demonstrate that compared to traditional RL, Safe RL, Offline RL, IL, and TSF algorithms, the S2CD framework significantly enhances both learning efficiency and model performance, while also reducing training costs. Most importantly, even with suboptimal teacher performance, the S2CD framework enhances safety during training. The framework is also applicable beyond varying simulation environments and offers a foundation for future knowledge transfer research between simulation and real-world environments.
The primary limitation of this study is that all experiments were conducted in simulated environments, without utilizing teacher models trained in simulation environments to guide student models in real-world conditions. Therefore, the next step in our research will involve real-world vehicle experiments to further validate the effectiveness and practicality of the proposed framework. Additionally, this study did not consider mixed traffic scenarios involving both autonomous and non-autonomous vehicles, which represents another crucial direction for future research to expand the applicability and coverage of the S2CD framework.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (NSFC) under grant number 51975194.
@@ Appendix A Proof of Theorem 3
Using the mixed behavior policy as defined in Eq. 19, the difference between $J_{\theta_{\text{mix}}}$ and the return of the teacher’s policy $J_{\theta_{t}}$, is expressed as follows:
@@ Appendix C Hyper-parameters
Total steps of one episode 5,000
@@
Total steps of one episode 5,000
@@
Total steps of one episode 5,000
@@
Total steps of one episode 5,000
@@
Total steps of one episode 5,000
@@
Total steps of one episode 5,000

### v4

@@ 6 Conclusion
This paper introduces a novel framework, S2CD, based on knowledge transfer techniques. The S2CD framework first trains a teacher model in a simplified simulation environment and then utilizes this model to guide the student agent in a more complex environment, enhancing the safety and efficiency of the training process. The paper also presents an innovative RL algorithm, ACPPO+, which leverages samples generated by both teacher and student policies while dynamically adjusting the clipping factor based on sample importance, thereby improving learning efficiency. Moreover, the KL divergence between the teacher’s and student’s policies is incorporated as a constraint in model updates, solved using the Lagrangian method to enable the student agent to quickly adopt the policy of teacher. To enhance the student agent’s performance, a gradual weaning strategy reduces teacher intervention, ensuring that the student becomes more autonomous and independently explores optimal policies. Experimental results in highway lane-change scenarios demonstrate that compared to traditional RL, IL, Safe RL, TSF, and Offline RL, the S2CD framework significantly enhances both learning efficiency and model performance, while also reducing training costs. Most importantly, even with suboptimal teacher performance, the S2CD framework enhances safety during training. The framework is also applicable beyond varying simulation environments and offers a foundation for future knowledge transfer research between simulation and real-world environments.
The primary limitation of this study is that all experiments were conducted in simulated environments, without utilizing teacher models trained in simulation environments to guide student models in real-world conditions. Therefore, the next step in our research will involve real-world vehicle experiments to further validate the effectiveness and practicality of the proposed framework. Additionally, this study has not addressed mixed traffic scenarios involving intelligent and traditional vehicles, which represents an important direction for future research to expand the applicability and coverage of the S2CD framework.
@@ Appendix A Proof of Theorem 3
Using the mixed behavior policy as defined in Eq. 16, the difference between $J(\theta_{\text{mix}})$ and the return of the teacher’s policy $J(\theta_{t})$, is expressed as follows:
@@ Appendix C Hyper-parameters
Total steps per episode 5,000
@@
Total steps per episode 5,000
@@
Total steps per episode 5,000
@@
Total steps per episode 5,000
@@
Total steps per episode 5,000
@@
Total steps per episode 5,000